# Assessing treatment effects in observational data with missing confounders: A comparative study of practical doubly-robust and traditional missing data methods

### v1

@@ 2.2 Estimands
A fourth estimand of interest is the odds ratio conditional on covariates (we will often use the shorthand conditional odds ratio or cOR). This is the natural parameter from a logistic regression model, and is the target of interest in many applied settings. Suppose that we fit the following logistic regression model to the ideal data (with no missingness): $\operatorname{logit}P(Y=1\mid X=x,Z=z,W=w)=\beta_{0}+\beta_{1}x+\beta_{2}z+\beta_{3}w$. In this model, $\exp(\beta_{1})$ is the cOR comparing $X=1$ to $X=0$ conditional on Z and W; we will consider this to be the cOR of interest. It is conditional because the interpretation of $\beta_{1}$ is the log odds for a fixed level of $Z$ and $W$. While we focus on the cOR, the conditional risk difference (cRD) and conditional relative risk (cRR) could easily be estimated using a linear or Poisson working regression model, respectively.
@@
This distinction allows us to define two levels of estimand. We define the oracle estimand as a function of only the data-generating distribution (typically unknown); the mRD, mRR, and mOR as defined above are all examples of oracle estimands. If the outcomes are generated from the same model as the working logistic regression model, then cOR is also an oracle estimand. In contrast, we define the census estimand as a function of both the data-generating distribution and a working regression model. We can define a census version of each estimand listed above, using the working regression model in place of the true conditional probability of outcome given the covariates. For the marginal parameters defined above, the census version replaces the true treatment-specific mean outcome values with expectations of the outcome probabilities under the working model. The census estimand represents the natural target of estimation for certain methods, as we will highlight below, and performance for estimating the census estimand is often considered when comparing estimators in missing-data contexts. Importantly, if the working regression model corresponds to the data-generating model, then the oracle and census estimands are identical. If the goal of the analysis is to understand the true treatment effect, then the oracle estimand should be the target of estimation and a method whose natural target is the census estimand will be biased unless the census and oracle estimands are identical. Conversely, if the goal of the analysis is to estimate the association in a posited multivariable regression model, then the census estimand is the target of inference.
@@ 2.3 Missing data analysis approaches
Our goal is to estimate and make inference on the estimands defined above, using $n$ independent and identically distributed observations $O_{1},\ldots,O_{n}$. We will compare several estimators that handle the missing confounder data, including model-based estimators (including MI), design-based estimators (IPW and GR), and nonparametric estimators (TMLE). Our goal is to compare approaches that are readily implemented in current software, including the approaches common in practice. To this end, we define the missing-data model $\pi(y,x,z)=P(R=1\mid Y=y,X=x,Z=z)$, the propensity score model $g(z,w)=P(X=1\mid Z=z,W=w)$, and the outcome model $Q(x,z,w)=E(Y\mid X=x,Z=z,W=w)$. For simplicity, we will often drop the arguments for these functions and refer to them as $\pi$, $g$, and $Q$.
@@
We consider MI by chained equations (MICE; Van Buuren, 2007; Raghunathan et al., 2001) as the primary MI approach, due to its popularity in applied settings and its relative success compared to other imputation approaches in the setting of missing data in the confounders (Weberpals et al., 2024). Briefly, MICE requires choosing a number of imputations $M$, a number of iterations $K$ (controls the convergence of each imputation run), and an imputation method for each variable with missing data. Rather than relying on a joint distribution for imputations, at each iteration, MICE imputes each $\{W_{j}\}_{j=1}^{p}$ conditional on the other variables in a specified sequence (i.e. chain) of equations (Austin et al., 2021). After the $K$ iterations, this results in $M$ imputed datasets $O_{n,m}=(Y,X,Z,W_{m})$. We implemented MICE using the R package mice (van Buuren and Groothuis-Oudshoorn, 2011), choosing 20 imputations, a maximum of 25 iterations, and the default imputation method for each variable (van Buuren and Groothuis-Oudshoorn, 2011). We also considered multiple imputation with random forest (MI-RF), which also did well in Weberpals et al. (2024), and with XGBoost (Chen et al., 2015) (MI-XGB), anticipating potential advantages using this flexible method for our most complex simulation settings. MI-RF was implemented using the mice package and default settings. MI-XGB was implemented using the mixgb package (Deng, 2023), with five-fold cross-validation to select the optimal number of boosting rounds. In all cases, after performing MI, we fit a logistic regression model to estimate the outcome probability on each imputed dataset $O_{n,m}$ and pooled the results, using Rubin’s rules for inference (Rubin, 1987).
@@
We consider two design-based estimators: IPW and GR. Both estimators are based on weighted complete-case estimating equations. For both, we first fit a logistic regression model to estimate the probability of observing data; the estimated weights $\hat{\pi}_{i,\text{IPW}}^{-1}=\widehat{P}(R=1\mid Y_{i},X_{i},Z_{i})^{-1}$ resulting from this logistic regression model fit are the final weights for IPW. Survey calibration estimators use auxiliary variables available on the entire cohort to adjust the design weights, here estimated by $\hat{\pi}_{i,\text{IPW}}^{-1}$, in a way that leverages information in the auxiliary data regarding the target of estimation. When the target of inference is one or more regression coefficients, the random vector of the expected value of the efficient influence functions (EIFs, Bickel et al., 1993) for the regression coefficients given the observed phase I data (in our setting, given the variables derived from claims data), denoted $h_{i}$, is the ideal auxiliary variable for calibration which estimates the design-efficient AIPW estimator (Breslow et al., 2009a; Lumley et al., 2011). Specifically, for GR, we obtain calibrated weights $\hat{\pi}_{i,\text{GR}}^{-1}=a_{i}\hat{\pi}^{-1}_{i,\text{IPW}}$, defined by the solution to
$\min_{a\in\mathbb{R}^{n}}$ $\sum_{i=1}^{n}R_{i}d(a_{i}\hat{\pi}_{i,\text{IPW}}^{-1},\hat{\pi}_{i,\text{IPW}}^{-1})$
subject to $\sum_{i=1}^{n}\hat{h}_{i}=\sum_{i=1}^{n}R_{i}a_{i}\hat{h}_{i}\hat{\pi}_{i,\text{IPW}}^{-1},$
@@
The MI, IPW, and GR strategies all naturally target the census cOR. They also provide an estimator of the outcome probability, which we can use to estimate marginal quantities. Suppose that we have an estimator $Q_{n}(x,z,w)=\widehat{P}(Y=1\mid X=x,Z=z,W=w)$ obtained using logistic regression. Suppose further that we have a set of weights $\hat{\pi}_{i}$; for MI, $\hat{\pi}_{i}=1$ for $i=1,\ldots,n$, while for IPW and GR $\hat{\pi}_{i}$ are (possibly calibrated) inverse probability weights. To estimate $\mu_{1}$ (the mean outcome value under treatment), we can first use standard software to predict using the fitted logistic regression model after setting $X=1$ for all observations, yielding $Q_{n}(1,Z_{i},W_{i})=\widehat{P}(Y=1\mid X_{i}=1,Z_{i},W_{i})$; we then take an average over the empirical distribution of the covariates:
$\mu_{1,n}=\frac{1}{n}\sum_{i=1}^{n}\hat{\pi}_{i}Q_{n}(1,Z_{i},W_{i}).$
The same procedure applied after setting $X=0$ for all observations yields an estimator $\mu_{0,n}$ of $\mu_{0}$. These estimators can be plugged in to yield estimators of the mRD, mRR, and mOR:
@@
We obtain the fluctuation using logistic regression with outcome $Y$, covariates $[H_{1}(X_{i},Z_{i},W_{i}),H_{0}(X_{i},\allowbreak Z_{i},W_{i})]$, offset $Q_{n}(X_{i},Z_{i},W_{i})$, and weights
@@
Analysis approach Required models for estimation Correct specification for consistent estimation of oracle parameters Correct specification for consistent estimation of census parameters
IPW $Q$ and $\pi$ $Q$ and $\pi$ $\pi$ or (Q and CD-MCAR)
MI $Q$ and $f$ $Q$ and $f$ $f$
GR $Q$ and $\pi$ $Q$ or $\pi$ $\pi$ or $Q$
IPCW-TMLE $Q$, $g$ and $\pi$ ($Q$ or $g$) and $\pi$ $\pi$ or [($Q$ or $g$) and CD-MCAR]
Table 1: Model specifications for missing-data estimation procedures and requirements for consistency for census and oracle estimands, under missing-at-random and assuming that there are no unmeasured confounders. The working outcome model $Q$ is assumed to be the same as the census model. The missing-data model is $\pi$, the imputation model is $f$, and the treatment/exposure propensity score model is $g$.
  • IPW: inverse probability weighted outcome regression where weights account for missing data; CD-MCAR: covariate-dependent missing completely at random, where missing data can depend only on always-observed covariates, not outcomes (Seaman et al., 2013); MI: multiple imputation; GR: generalized raking; IPCW: inverse probability of coarsening weighted; TMLE: targeted minimum loss-based estimation

### v2

@@ 2.2 Estimands
A fourth estimand of interest is the odds ratio conditional on covariates (we will often use the shorthand conditional odds ratio or cOR). This is the natural parameter from a logistic regression model, and is the target of interest in many applied settings. This is the natural parameter from a logistic regression model, and is the target of interest in many applied settings. Suppose that we fit the following logistic regression model to the ideal data (with no missingness): $\operatorname{logit}P(Y=1\mid X=x,Z=z,W=w)=\beta_{0}+\beta_{1}x+\beta_{2}z+\beta_{3}w$. In this model, $\exp(\beta_{1})$ is the cOR comparing $X=1$ to $X=0$ conditional on Z and W; we will consider this to be the cOR of interest. It is conditional because the interpretation of $\beta_{1}$ is the log odds for a fixed level of $Z$ and $W$. While we focus on the cOR, the conditional risk difference (cRD) and conditional relative risk (cRR) could easily be estimated using a linear or Poisson working regression model, respectively.
@@
This distinction allows us to define two levels of estimand. We define the oracle estimand as a function of only the data-generating distribution (typically unknown); the mRD, mRR, and mOR as defined above are all examples of oracle estimands. If the outcomes are generated from the same model as the working logistic regression model, then the cOR is also an oracle estimand. In contrast, we define the census estimand as a function of both the data-generating distribution and a working regression model. We can define a census version of each estimand listed above, using the working regression model in place of the true conditional probability of outcome given the covariates. For the marginal parameters defined above, the census version replaces the true treatment-specific mean outcome values with expectations of the outcome probabilities under the working model. The census estimand represents the natural target of estimation for certain methods, as we will highlight below, and performance for estimating the census estimand is often considered when comparing estimators in missing-data contexts. Importantly, if the working regression model corresponds to the data-generating model, then the oracle and census estimands are identical. If the goal of the analysis is to understand the true treatment effect, then the oracle estimand should be the target of estimation and a method whose natural target is the census estimand will be biased unless the census and oracle estimands are identical. Conversely, if the goal of the analysis is to estimate the association in a posited multivariable regression model, then the census estimand is the target of inference.
@@ 2.3 Missing data analysis approaches
Our goal is to estimate and make inference on the estimands defined above, using $n$ independent and identically distributed observations $O_{1},\ldots,O_{n}$. We will compare several estimators that handle the missing confounder data, including model-based estimators (including MI), design-based estimators (IPW and GR), and nonparametric estimators (TMLE). Our goal is to compare approaches that are readily implemented in current software, including the approaches common in practice. To simplify the notation, for the remainder of this manuscript we consider a setting where there is a single indicator $R$ of whether all of $W$ is observed. To this end, we define the missing-data model $\pi(y,x,z)=P(R=1\mid Y=y,X=x,Z=z)$, the propensity score model $g(z,w)=P(X=1\mid Z=z,W=w)$, and the outcome model $Q(x,z,w)=E(Y\mid X=x,Z=z,W=w)$. For simplicity, we will often drop the arguments for these functions and refer to them as $\pi$, $g$, and $Q$.
@@
We consider MI by chained equations (MICE; Van Buuren, 2007; Raghunathan et al., 2001) as the primary MI approach, due to its popularity in applied settings and its relative success compared to other imputation approaches in the setting of missing data in the confounders (Weberpals et al., 2024). Briefly, MICE requires choosing a number of imputations $M$, a number of iterations $K$ (controls the convergence of each imputation run), and an imputation method for each variable with missing data. Rather than relying on a joint distribution for imputations, at each iteration, MICE imputes each $\{W_{j}\}_{j=1}^{q}$ conditional on the other variables in a specified sequence (i.e. chain) of equations (Austin et al., 2021). After the $K$ iterations, this results in $M$ imputed datasets $O_{n,m}=(Y,X,Z,W_{m})$. We implemented MICE using the R package mice (van Buuren and Groothuis-Oudshoorn, 2011), choosing 20 imputations, a maximum of 25 iterations, and the default imputation method for each variable (van Buuren and Groothuis-Oudshoorn, 2011). We also considered multiple imputation with random forest (MI-RF), which also did well in Weberpals et al. (2024), and with XGBoost (Chen et al., 2015) (MI-XGB), anticipating potential advantages using this flexible method for our most complex simulation settings. MI-RF was implemented using the mice package and default settings. MI-XGB was implemented using the mixgb package (Deng, 2023), with five-fold cross-validation to select the optimal number of boosting rounds. In all cases, after performing MI, we fit a logistic regression model to estimate the outcome probability on each imputed dataset $O_{n,m}$ and pooled the results, using Rubin’s rules for inference (Rubin, 1987).
@@
We consider two design-based estimators: IPW and GR. Both estimators are based on weighted complete-case estimating equations. For both, we first fit a logistic regression model to estimate the probability of observing data; the estimated weights $\hat{\pi}_{i,\text{IPW}}^{-1}=\widehat{P}(R=1\mid Y_{i},X_{i},Z_{i})^{-1}$ resulting from this logistic regression model fit are the final weights for IPW. Survey calibration estimators use auxiliary variables available on the entire cohort to adjust the design weights, here estimated by $\hat{\pi}_{i,\text{IPW}}^{-1}$, in a way that leverages information in the auxiliary data regarding the target of estimation. When the target of inference is one or more regression coefficients, the random vector of the expected value of the efficient influence functions (EIFs, Bickel et al., 1993) for the regression coefficients given the observed phase I data (in our setting, given the variables derived from claims data), denoted $h_{i}$, is the ideal auxiliary variable for calibration which estimates the design-efficient AIPW estimator (Breslow et al., 2009a; Lumley et al., 2011). Specifically, for GR, we obtain calibrated weights $\hat{\pi}_{i,\text{GR}}^{-1}=a_{i}\hat{\pi}^{-1}_{i,\text{IPW}}$, where $a$ is defined by
$a=\operatorname{argmin}_{a^{\prime}\in\mathbb{R}^{n}}$ $\sum_{i=1}^{n}R_{i}d(a_{i}^{\prime}\hat{\pi}_{i,\text{IPW}}^{-1},\hat{\pi}_{i,\text{IPW}}^{-1})$
subject to $\sum_{i=1}^{n}\hat{h}_{i}=\sum_{i=1}^{n}R_{i}a_{i}^{\prime}\hat{h}_{i}\hat{\pi}_{i,\text{IPW}}^{-1},$
@@
The MI, IPW, and GR strategies all naturally target the census cOR, because they are based on the working parametric outcome regression model. They also provide an estimator of the outcome probability, which we can use to estimate marginal quantities. Suppose that we have an estimator $Q_{n}(x,z,w)=\widehat{P}(Y=1\mid X=x,Z=z,W=w)$ obtained using logistic regression. Suppose further that we have a set of weights $\hat{\pi}_{i}^{-1}$; for MI, $\hat{\pi}_{i}^{-1}=1$ for $i=1,\ldots,n$, while for IPW and GR $\hat{\pi}_{i}^{-1}$ are inverse probability weights (calibrated weights for GR). To estimate $\mu_{1}$ (the mean outcome value under treatment), we can first use standard software to predict using the fitted logistic regression model after setting $X=1$ for all observations, yielding $Q_{n}(1,Z_{i},W_{i})=\widehat{P}(Y=1\mid X_{i}=1,Z_{i},W_{i})$; we then take an average over the empirical distribution of the covariates:
$\mu_{1,n}=\frac{1}{n}\sum_{i=1}^{n}R_{i}\hat{\pi}^{-1}_{i}Q_{n}(1,Z_{i},W_{i}).$
This procedure requires that $\hat{\pi}$ is a consistent estimator of $\pi$ and that $Q_{n}$ is a consistent estimator of $Q$. We do not implement further augmentation (Breslow et al., 2009a) because our aim was to compare commonly-used approaches (without augmentation) with augmented procedures like GR. The same procedure applied after setting $X=0$ for all observations yields an estimator $\mu_{0,n}$ of $\mu_{0}$. These estimators can be plugged in to yield estimators of the mRD, mRR, and mOR:
@@
We obtain the fluctuation using logistic regression with outcome $Y$, covariates $H_{1}(X_{i},Z_{i},W_{i}),H_{0}(X_{i},\allowbreak Z_{i},W_{i})]$, offset $Q_{n}(X_{i},Z_{i},W_{i})$, and weights
@@ 3.1.6 Estimators
3.1.7 Performance metrics
3.2 Simulation results
3.2.1 Base case (12% outcome probability, 40% MAR missing)
Figure 1 (Figure 2) shows the estimator relative performance for the cOR census (oracle) estimand for the base case simulations (detailed results in Tables S6S13). In the simple outcome and MAR scenario, the oracle and census estimands are the same and all methods, except the CNFD and CC approaches, maintained 95% nominal coverage. The GR estimator is consistent for the census estimand in all MAR cases because the analytic working model is the census model and so, as expected, maintained correct coverage for this parameter for these settings; unexpectedly, GR also maintained close to the correct coverage for both census and oracle estimands when the analytic working models both were misspecified (i.e. under the complex outcome and MAR scenario). GR also maintained the best efficiency for the census estimand and generally within about 5% of the best efficiency for the oracle estimand (i.e. lowest rRMSE) relative to the other methods for all 4 combinations of simple versus complex data generating scenarios. Conversely, MICE suffered from bias and inefficiency when the true outcome model was complex (i.e., the working model was misspecified); this bias was worse for the census estimand. The MI procedures, even in the complex outcome scenario, were still more efficient than CC, IPW and the TMLE estimators; however, which imputation procedure did the best varied by setting. Generally the MI estimates were close when the outcome DGM was simple, with an apparent very slight advantage to MICE. When the outcome scenario was complex MI-RF had lower rRMSE than MICE for both the census and oracle estimands; MI-XGB tended to have lower rRMSE for the oracle cOR estimand while MI-RF had lower rRMSE for the census estimand (results for MI-XGB are only displayed in Tables S6S8, S10S12). The TMLE estimators were generally more efficient than IPW and CC, but not as efficient as MI and GR. Further, TMLE estimators, which were implemented as weighted two-phase estimators, suffered from some bias when the missing data model was more complex, with oracle coverage only around 92%; however, IPW fared much worse for the complex MAR model, with oracle (and nominal) coverage dropping to about 70%. TMLE-M was less efficient than TMLE-MTO for the complex outcome scenario. The CNFD method demonstrated severe bias, with nominal coverage under 20%, across all scenarios.
Figure 1: Synthetic Data MAR Simulation: Census clogOR. Comparing estimators of the census clogOR estimand for the base case with 40% missingness and 12% outcome probability, where the data generating models were simple or complex. Top graph: Percent Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True clogOR values are 0.405 and 0.371 for simple and complex outcome models, respectively.
Figure 2: Synthetic Data MAR Simulation: Oracle clogOR. Comparing estimators of the oracle clogOR estimand for the base case with 40% missingness and 12% outcome probability, where the data generating models were simple or complex. Top graph: Percent Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true clogOR value is 0.405 for both simple and complex models.
3.2.2 Scenarios with varying outcome and missing data probabilities
3.2.3 MNAR Scenarios
Figure 3: Synthetic Data MNAR Simulation: Census clogOR. Comparing estimators of the census estimand with 40% confounder missingness and 12% outcome probability. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True clogOR values are 0.405 for simple outcome models and 0.404 and 0.371 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Figure 4: Synthetic Data MNAR Simulation: Oracle clogOR. Comparing estimators of the oracle estimand with 40% missingness and 12% outcome probability. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The correct clogOR value is 0.405 across all scenarios.
4.1 Study cohort
4.2 Plasmode simulation methods
We seek to compare the same estimation methods described in Section 3.1.6 for four sets of plasmode simulation studies, where we vary the rarity of the outcome and the complexity of the underlying DGMs. We used two different DGMs that were derived by fitting different models to the KPWA cohort: 1) a parametric generalized linear model (GLM) (i.e., logistic regression) with a few interactions and 2) a tree-based approach allowing for complex interactions. These two DGMs allow us to investigate the performance of estimators under a simple and more complex DGM, where the latter was chosen to be have a functional form notably different than the typical linear parametric working model. In both DGMs, we used the same approach (i.e., GLM or trees) to estimate the propensity score, outcome regression, and missing-data models. We only consider the BNMK-O estimator for the GLM DGM because there is no single oracle cOR of interest in the tree-based DGM.
We began by estimating the missing-data model using the entire sample of 112,770 individuals; in this model, the outcome was an indicator of whether the PHQ was observed or not. Next, we estimated the treatment propensity score and outcome regression models using the complete-data cohort (N=50,337). Importantly, using this complete-data cohort is convenient for plasmode simulation, but precludes making general conclusions about the larger population. Table S23 and Figures S36S38 describe the final parametric and tree-based models, respectively. With these models fixed, we generate plasmode data as follows. We use bootstrap sampling with replacement from the complete-data cohort (i.e., individuals with fully observed PHQ-9) to generate complete covariate data for each person for 1000 plasmode datasets of size N=50,337, thereby maintaining the complex data structure in the KPWA cohort. For each DGM approach, we use the fitted models to generate treatment given the plasmode sample of covariates, the outcome given the treatment and covariates, and then the missing data indicators for the PHQ-9 given the treatment, covariates and outcome for the 1000 plasmode datasets. Importantly, the data are MAR because the probability of missing data depends only on observed variables. In all cases, the working models that we use for analysis are misspecified, because the true data-generating models contain interactions that are not modeled (Table S23). The plasmode sampling for the simulations drew only from individuals with PHQ-9 measured (i.e., the complete-data cohort of size 50,337) and we use the generated missing data indicators to “hide” PHQ information for selected individuals.
Similar performance metrics as for the synthetic data simulations were calculated for each estimator across the plasmode simulations. Because the value of the DGM treatment effect estimands derived from the KPWA were close to null, bias instead of percent bias was calculated. The estimators returned results in >98>98% of plasmode datasets (Section S3.4).
4.3 Plasmode simulation results
Figure 5: Plasmode Data Simulation: Census mRD. Estimator relative performance for the census mRD estimand, across GLM and tree-based data generating models for the rare and common outcomes. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$ True values of mRD from lightest to darkest scenarios are 7.6E-4, 1.2E-4, -1.6E-2, and -6.2E-3.
Figure 6: Plasmode Data Simulation: Oracle mRD. Estimator relative performance for the oracle mRD estimand, across GLM and tree-based data generating models for the rare and common outcomes. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$ True values of mRD from lightest to darkest scenarios are 7.5E-4, -2.4E-4, -1.7E-2, and -7.6E-3.
5 Discussion
Our numerical experiments highlight the importance of first choosing a target estimand and then determining an estimation procedure. As expected, in cases where the working outcome regression model coincided with the data-generating outcome model and a simple MAR missing data model (the “base case” simulations), all methods performed well, with bias near zero and confidence interval coverage near the nominal 95% level. However, even in this standard case we noted several important findings. The doubly-robust GR approach, which reduces to a tailored re-weighting of complete cases, maintained nearly the same efficiency as MI, the method generally expected to be the most efficient for correctly-specified parametric models. While MI can be highly efficient, the MI method (e.g., MICE or MI-RF) that performs best can vary across scenarios, but we do not know the DGM scenario in practice; for this reason, GR is often a safer choice than MI. Perhaps less surprising, both the CNFD and CC approaches were biased and highly inefficient; and IPW and TMLE were inefficient. Importantly, the CNFD model — resulting from omitting the confounders with missing values — is always biased and is not recommended; all missing data approaches that we evaluated outperformed the CNFD approach.
Correct specification for consistent estimation of:
Analysis approach Required models for estimation oracle parameters census parameters in regression models marginal census parameters
IPW $Q$ and $\pi$ $Q$ and $\pi$ $\pi$ or CD-MCAR $\pi$ or (Q and CD-MCAR)
MI $Q$ and $f$ $Q$ and $f$ $f$ $f$
GR $Q$, $\pi$ and ($\eta$ or $\eta_{Q}$) $Q$ and ($\pi$ or $\eta$) $\pi$ or $\eta_{Q}$ or CD-MCAR ($\pi$ or $\eta_{Q}$) or ($Q$ and CD-MCAR)
IPCW-TMLE $Q$, $g$ and $\pi$ ($Q$ or $g$) and $\pi$ $\pi$ or [($Q$ or $g$) and CD-MCAR] $\pi$ or [($Q$ or $g$) and CD-MCAR]
Table 1: Model specifications for missing-data estimation procedures and requirements for consistency for census and oracle estimands, under missing-at-random and assuming that there are no unmeasured confounders. The working outcome model $Q$ is assumed to be the same as the census model. The missing-data model is $\pi$, the imputation model is $f$, the treatment/exposure propensity score model is $g$, and the augmentation term/optimal raking variable for the oracle parameter is $\eta$ (a property of the data-generating mechanism) and for the census parameter is $\eta_{Q}$ (which depends on both $Q$ and the data-generating mechanism).
  • IPW: inverse probability weighted outcome regression where weights account for missing data; CD-MCAR: covariate-dependent missing completely at random, where missing data can depend only on always-observed covariates, not outcomes (Seaman et al., 2013); MI: multiple imputation; GR: generalized raking; IPCW: inverse probability of coarsening weighted; TMLE: targeted minimum loss-based estimation.